### v1

# Likelihood-free inference of phylogenetic tree posterior distributions

Abstract
Phylogenetic inference, the task of reconstructing how related sequences evolved from common ancestors, is a central task in evolutionary genomics. The current state-of-the-art methods exploit probabilistic models of sequence evolution along phylogenetic trees, by searching for the tree maximizing the likelihood of observed sequences, or by estimating the posterior of the tree given the sequences in a Bayesian framework. Both approaches typically require to compute likelihoods, which is only feasible under simplifying assumptions such as independence of the evolution at the different positions of the sequence, and even then remains a costly operation. Here we present Phyloformer 2, the first likelihood-free inference method for posterior distributions over phylogenies. Phyloformer 2 exploits a novel encoding for pairs of sequences that makes it more scalable than previous approaches, and a parameterized probability distribution factorized over a succession of subtree merges. The resulting network provides accurate estimates of the posterior distribution, and outperforms both state-of-the-art maximum likelihood methods and a previous likelihood-free method for point estimation. It opens the way to fast and accurate phylogenetic inference under realistic models of sequence evolution.
1 Introduction
The genomes of living species evolve over time through a process that involves mutations and selection. Reconstructing the evolutionary history of a set of contemporaneous sequences is a central task in genomics (Kapli et al., 2020): it is used to understand how extant species have evolved from common ancestors (Álvarez Carretero et al., 2022), how bacterial resistances to drugs have emerged and been disseminated (Aminov & Mackie, 2007), and how epidemics are spreading Hadfield et al. (2018). A key object in this endeavor is the phylogeny, a bifurcating tree summarizing the succession of transformations of the sequence that lead to the current observed diversity from a single ancestor. For the past twenty years, the field of phylogenetic reconstruction has been largely dominated by approaches based on probabilistic models of sequence evolution (Lartillot & Philippe, 2004; Le & Gascuel, 2008; Muñoz-Gómez et al., 2022). These models are typically continous time markov processes parameterized by the phylogeny and other values such as the rates at which a given aminoacid or nucleotide substitutes into another one. Existing reconstruction methods look for the phylogeny maximizing the likelihood of the sequences under such a model (Minh et al., 2020; Price et al., 2010) or, in a Bayesian perspective, aim at sampling from the posterior distribution of the phylogeny given the sequences through Monte Carlo strategies (Huelsenbeck & Ronquist, 2001; Höhna et al., 2016; Bouchard-Côté et al., 2012) or approximate this posterior distribution through variational inference (Zhang & IV, 2019; Xie & Zhang, 2023; Zhou et al., 2024; Duan et al., 2024). Across all these approaches, a major hurdle is the computational cost of evaluating the likelihood function which is required for numerical optimization in maximum likelihood estimators, and to compute acceptance probability in sampling strategies or the evidence lower bound (ELBO) objective to be maximized in variational inference. The likelihood for any single phylogeny is computed through a costly pruning algorithm (Felsenstein, 1981) and exploring the set of possible tree topologies—(2n5)!!(2n-5)!! for a phylogeny over $n$ leaves—even heuristically requires many such evaluations. Furthermore, just making this computation feasible has forced the evolutionary genomics community to focus its effort on probabilistic models that make simplifying assumptions such as independence and identical distribution of the evolution at each position in the sequence, or the absence of natural selection. These simplifications are known to produce unrealistic sets of sequences (Trost et al., 2024), artifacts in reconstructed phylogenies (Telford et al., 2005), and imped our ability to understand the evolutionary history of living species.
Simulation-based or likelihood-free inference has emerged as a powerful paradigm for estimation under probabilistic models under which likelihood evaluations are intractable but sampling is cheap (Cranmer et al., 2020; Lueckmann et al., 2021). In particular, this paradigm has leveraged advances in deep learning to produce methods that approximate posterior distributions by neural networks trained over data simulated under probabilistic models (Greenberg et al., 2019; Lueckmann et al., 2021). Among these methods, neural posterior estimation (NPE, Lueckmann et al., 2021) defines a family of distribution parameterized by a neural network whose weights are then optimized to approximate the posterior. In addition to working around the need for likelihood evaluations, NPE is amortized: training the network can take time but performing inferences with the trained network is typically very fast. However, defining a parameterized family of distributions that is appropriate for the posteriors of phylogenies given sequences is not straighforward.
Here we introduce Phyloformer 2, a NPE for phylogenetic reconstruction, with the following contributions:
  • We propose a parameterized family of posterior distributions on phylogenies given a set of sequences, factorized through a succession of pairwise mergings. Optimizing the weights of our network to maximize the log-probability within this family yields a likelihood-free estimate of the corresponding posterior that we can use to evaluate the probability of a phylogeny, for sampling, or to produce the maximum a posteriori. To our knowledge, this is the first likelihood-free posterior estimation method trained end-to-end from sequences to the phylogeny beyond quartets.
  • To extract the parameters of the approximate posterior distribution from the input sequences, we introduce a novel architecture akin to the EvoFormer module used in Alphafold 2, that is both more scalable and expressive than the one used in Phyloformer. The overall architecture allows us to scale up to over 200 sequences of length 500 or more than 300 sequences of length 250 on a single V100 GPU with 16Gb of VRAM.
  • On data generated under a probabilistic model of sequence evolution with tractable likelihood, Phyloformer 2 outperforms both state-of-the-art likelihood-based and likelihood-free reconstruction methods in topological accuracy and produces estimates of the posterior compared to MCMC samples. Because it is likelihood-free, it can also be trained to produce estimates under models with intractable likelihoods, in which case the performance gap with—misspecified—likelihood-based estimators further increases.
  • Because Phyloformer 2 is amortized, once trained, it can perform inference 1 to 2 orders of magnitude faster than the—less accurate—state-of-the-art likelihood-based estimators.
Related work
Initial attempts to phylogenetic NPE have restricted themselves to quartets, i.e., topologies over four leaves, allowing them to cast the problem as a classification over the three possible topologies (Suvorov et al., 2019; Zou et al., 2020; Tang et al., 2024). In this case, a vector embedding of the input sequences was extracted by a neural network and used to produce three scalar outputs, and the probability of each topology was simply modeled as a softmax over these three outputs. Because the number of possible topologies grows super-exponentially with the number of leaves, this strategy cannot be generalized to larger numbers of sequences and even if it could, treating all topologies as separate classes would disregard the fact that some are more similar than others. In addition, Grosshauser M (2021) re-evaluated the method of Zou et al. (2020) and showed that it underperformed on more difficult tasks with short sequences and long evolution times. Alternatively, Nesterenko et al. (2025) proposed Phyloformer, a network predicting evolutionary distances, i.e., sum of branch lengths on the phylogeny between pairs of leaves, minimizing the mean absolute error (MAE) between true and predicted distances. Given all correct pairwise distances, an existing algorithm (neighbor joining, NJ, Saitou & Nei, 1987) is guaranteed to reconstruct the correct tree, but the authors observed limited topological reconstruction accuracy suggesting an irreducible discrepancy between the two metrics, increasingly so for larger numbers of leaves. In addition, minimizing the MAE against scalar values produced by the network only allows for the estimation of a point estimate—namely the median of the posterior rather than the entire distribution. Finally, the network applied axial self-attention (Ho et al., 2019) to all pairs of sequences, which led to a large memory footprint even using a linear approximation of self-attention (Katharopoulos et al., 2020) and limiting the scalability of the approach to 200 sequences of length 500 in practice. The authors reported that using axial self-attention on sequences instead of pairs as in the MSA transformer (Rao et al., 2021) improved the scalability but dramatically decreased the reconstruction accuracy.
2 Background and notation
2.1 Notation
Let $x=\{x_{1},\ldots,x_{N}\}$ be a set of $N$ sequences of $L$ letters in some fixed alphabet representing organic compounds (e.g., 20 possible amino acids for proteins, 4 possible nucleotides for DNA). We assume that the sequences are aligned, i.e., correspond to $N$ initial sequences of possibly different lengths whose positions were matched to minimize some score quantifying how similar all sequences are at each position, potentially by introducing gaps denoted by a special character in the alphabet (Kapli et al., 2020). A phylogeny $\theta=(\tau,\ell)$ over $x$ is an unrooted binary tree $\tau$ with $N$ leaves and a set $\ell$ of branch lengths in $\mathbb{R}^{2N-3}_{+}$. $\tau$ can equivalently be represented by a succession of merges. Intuitively, given N species, one chooses two species, pair them to form a cherry, replace them by a single species representing their ancestor, and proceed recursively with the N-1 resulting species until the entire tree has been produced. Formally, we denote this succession of merges as $\left\{m^{(k)}\right\}_{k=1}^{N-3}$ where $m^{(k)}\in\left\{\left(v^{(k)}_{i},v^{(k)}_{j}\right)\in\mathcal{S}_{(k)}^{2}|i\neq j\right\}$ is a pair of two distincts elements in $\mathcal{S}_{(k)}$ the set of “mergeable” nodes at the $k^{th}$ merge—i.e., either leaves or internal nodes whose children were already merged. $\mathcal{S}_{(1)}$ is the set of leaves in $\tau$ and for $k>1$, $\mathcal{S}_{(k)}\stackrel{{\scriptstyle\Delta}}{{=}}\left\{\left\{\mathcal{S}_{(k-1)}\cup u^{(k-1)}\right\}\setminus\left(v^{(k-1)}_{i},v^{(k-1)}_{j}\right)\right\}$, and $u^{(k)}$ is the common neighbor in $\tau$ to the nodes merged in $m^{(k)}$—which is always defined because we start from leaves and recursively replace pairs of elements by their common neighbor in the binary tree. We further denote $\ell^{(k)}=\left\{\ell^{(k)}_{i},\ell^{(k)}_{j}\right\}\in\mathbb{R}_{+}^{2}$ the $k$-th cherry, i.e., the set of two branch lengths connecting $m^{(k)}$ to $u^{(k)}$ . In our context, the $N$ leaves will represent the taxa with sequences in $x$, $u^{(k)}$ is the common ancestor of $m^{(k)}$, and $\ell^{(k)}$ is the set of evolutionary times between this ancestor and its two descendants.
2.2 Neural posterior estimation
For a probabilistic model $p(x|\theta)$ of the data $x$ given the parameter $\theta$ and a prior distribution $p(\theta)$, NPE provides a way to estimate the posterior $p(\theta|x)$ in cases where evaluating $p(x|\theta)$ for a given $(x,\theta)$ is too costly or intractable, but where it is possible to sample from this model. It relies on a family of distributions $q_{\psi}(\theta|x)$ whose parameters $\psi$ are provided by a neural network acting on the data xx—by abuse of notation we denote $\psi(x)$ the parameters output by this neural network for an input $x$. NPE builds its approximation of $p(\theta|x)$ by looking for the $q_{\psi}$ minimizing the Kullback-Leibler (KL) divergence $\mathbb{E}_{p(x)}\left[\text{KL}(q_{\psi}(\theta|x)||p(\theta|x))\right]$ with the true posterior. This is generally achieved by minimizing $\sum_{i=1}^{n}\log q_{\psi(x_{i})}(\theta_{i}|x_{i})$ over a large number of examples $\{(x_{i},\theta_{i})\}_{i=1}^{n}$ sampled from $p(x,\theta)$ by successively sampling a $\theta_{i}$ from the prior and an $x_{i}$ from the model given $\theta_{i}$. These samples are therefore used to build a Monte Carlo approximation of $\mathbb{E}_{p(x,\theta)}\left[\log q_{\psi(x)}(\theta|x)\right]$, whose maximization is equivalent to minimizing the target average KL divergence (see e.g., Radev et al., 2020). Consequently, NPE is guaranteed to converge to the true posterior $p(\theta|x)$, provided that the family $q_{\psi}$ is expressive enough to represent $p(\theta|x)$. More precisely, its approximation error depends both on the expressivity of the chosen family of distributions, i.e., on the average KL divergence $\mathbb{E}_{p(x)}\left[\text{KL}(q_{\psi^{*}(x)}(\theta|x)||p(\theta|x))\right]$ between the best distribution $q_{\psi^{*}(x)}$ in the family, and the true posterior for each $x$, and on the expressivity of the neural network, i.e., on its ability to map every $x$ to the corresponding best parameters $\psi^{*}(x)$.
3 Methods
Phyloformer 2 combines two novel modules. The first one, EvoPF, encodes distribution parameters $\psi(x)$ from a set of aligned related sequences $x$, which is sometimes refered to as a multiple sequence alignment (MSA). The second one, BayesNJ, defines a family of posterior distributions $q_{\psi(x)}\left(\theta=(\tau,\ell)|x\right)$ on phylogenetic trees, parameterized by the output of EvoPF. Figure 1a-b represents the overall architecture.
Figure 1: Architecture of Phyloformer 2. Panel a: EvoPF, an EvoFormer-inspired module updating $N\times L$ embeddings for a set of aligned sequences (MSA) $x$ and N(N1)/2N(N-1)/2 embeddings for pairs of sequences. Each of its blocks applies self-attention within both the MSA and representation, and ensures information sharing between them. After 12 blocks, we extract one embedding for each sequence by averaging the MSA embeddings across sites. Panel b: BayesNJ (Algorithm 1) computes the posterior probability of a phylogeny given an MSA represented by the sequence and pair embeddings provided by EvoPF. The probability is a product over a recursive operation where two taxa are merged into their parent, and the taxon representation is updated accordingly. Panel c: at inference time, we apply the same succession of operations as for evaluating the probability, but either sampling or taking the modes of the distributions (Algorithm S.10).
3.1 Encoding tree distribution parameters with EvoPF
Our encoder should process a set $x$ of aligned sequences and be expressive enough to capture sufficient information on their evolutionary relationships. Nesterenko et al. (2025) encoded $x$ with one embedding for each aligned position in each pair of aligned sequences. This approach avoided the need to flatten information across positions, while maintaining a representation at the pair level, consistently with their predicting pairwise evolutionary distances. They noted that tuning down the architecture to one embedding for each position in each sequence—boiling down to the MSA transformer (Rao et al., 2021)—dramatically improved the method scalability but strongly affected the accuracy of their trained network. Inspired by this trade-off, we introduce EvoPF, an encoder that maintains a single embedding per pair, and a separate representation with one embedding per position within each sequence (Figure 1a).
EvoPF is a transpose version of the EvoFormer module in Alphafold 2 Jumper et al. (2021)—whose objective was to capture spatial distances between pairs of aligned positions rather than evolutionary distances between pairs of homologous sequences—with a few simplifications. The MSA stack maintains an embedding for each position within each sequence in $x$. It relies on axial attention, alterning one layer of column-wise—i.e., between positions within each sequence separately—and row-wise self-attention—between sequences at each position separately. In parallel, we maintain an embedding for each pair of sequences, updated in each EvoPF block by flat self-attention between all pairs. This diverges from EvoFormer which relied on triangular attention where pair $(i,j)$ only attended to pairs $(i,k)$ and $(k,j)$. The MSA stack affects the pair stack through the addition of an outer product mean of sequence embeddings to the pair embeddings. Conversely, the pair stack affects the MSA stack by biasing its column-wise attention. We provide a complete description of EvoPF in Algorithms S.1 to S.9 in Section A.1.
3.2 Defining a proper probability distribution over phylogenies with BayesNJ
After 12 EvoPF blocks, we average the MSA embeddings across positions, yielding a single vector embedding per sequence. These embeddings constitute our encoding $\psi(x)$ of the sequences $x$, and our next goal is to define a family of distributions of phylogenies conditional on $x$, whose parameters are functions of $\psi(x)$. To this end, we define $q_{\psi(x)}\left(\theta=(\tau,\ell)|x\right)$ factorized over the succession of merges in $\tau$, i.e. $q_{\psi(x)}\left(\theta=(\tau,\ell)|x\right)=\Pi_{k=1}^{2N-3}q_{m}(m^{(k)}|m^{(<k)})q_{\ell}\left(\ell^{(k)}|m^{(k)},m^{(<k)}\right)$, where $m^{(<k)}$ denotes the set of merges with indices smaller than $k$. A caveat of this factorization is that most phylogenies $\theta$ can be obtained by several distinct successions of merges from the leaves. For example, a balanced binary tree with four leaves $a,b,c,d$ merging $(a,b)$ and $(c,d)$ can be obtained by merging any of the two groups first, and these two orders have no reason to lead to the same $\Pi_{k=1}^{2N-3}q_{m}(m^{(k)}|m^{(<k)})q_{\ell}\left(\ell^{(k)}|m^{(k)},m^{(<k)}\right)$ in general. Properly defining a probability distribution over phylogenies therefore requires to sum over all possible orders of merge, which is not feasible even for moderately large $N$.
Alternatively, we must ensure that for a given phylogeny our distribution assigns a non-zero probability to a single merge order and, to be able to evaluate the probability of any phylogeny, that this order can be recovered efficiently from the phylogeny. We achieve this by designing a canonical merge order such that (i) we can guarantee that our sampling procedure always generates merges in this order and (ii) our evaluation procedure always processes merges in this order. Point (i) requires that the merge order does not depend on the stochastic parts of the sampling procedure, point (ii) requires that we can recover the canonical order efficiently for any given phylogeny. We achieve these two points by ensuring that at every step $k$, we select the merge corresponding to the two closest nodes currently available in $\mathcal{S}_{(k)}$i.e., whose children have already been merged. Conversely when sampling a phylogeny $(\tau,\ell)$ from our $q_{\psi(x)}(\tau,\ell|x)$, we ensure that the distance between merged nodes $(i,j)$ is larger than the distance between any previous merges done while $(i,j)$ was a possible merge—i.e., after their children were merged.
3.3 BayesNJ parameterization of topological and branch-length components
We parameterize the topological component $q_{m}(m^{(k)}|m^{(<k)})$ by a softmin across pairwise scores computed from a bilinear function over embeddings of pairs of elements in $\mathcal{S}_{(k)}$, i.e., mergeable nodes at step $k$. At $k=1$ these embeddings are those provided by EvoPF, and at each successive merge we remove two current leaves and create an embedding for their ancestor which becomes a new leaf (Figure 1b). In order to ensure that each sampled merge $m^{(k)}$ leads to a longer cherry (sum of branch lengths) than those in all merges previously selected while $m^{(k)}$ was a possible choice, we re-parameterize the two branch length $(\ell^{(k)}_{i},\ell^{(k)}_{j})\in\ell^{(k)}$ as their sum $s^{(k)}=\ell^{(k)}_{i}+\ell^{(k)}_{j}$ and ratio $r^{(k)}=\ell^{(k)}_{i}/\ell^{(k)}_{j}$. We update a matrix of constraints as we merge pairs indicating the minimal length that the sum of branch lengths at any new merge must attain for the sampled tree to be consistent with our canonical order. We then model the probability $q_{s}(s^{(k)}|m^{(k)},m^{(<k)})$ as a Gamma distribution shifted by the constraint, i.e., $q_{s}(s^{(k)}|m^{(k)},m^{(<k)})=c_{m^{(k)}}+\text{Gamma}(\alpha,\lambda)$, where $c_{m^{(k)}}$ is the current constrain on the sum of branch length for merge $m^{(k)}$ and whose shape $\alpha$ and scale $\lambda$ are produced by a symmetric bilinear function of the embedding of the two merged nodes. We model the probability $q_{s}(r^{(k)}|m^{(k)},m^{(<k)})$ as a Beta distribution whose parameters are produced by a bilinear function of the embedding of the two merged nodes. We then obtain the joint probability $q_{\ell}\left(\ell^{(k)}|m^{(k)},m^{(<k)}\right)$ of the two branch lengths as $q_{s}(s^{(k)}|m^{(k)},m^{(<k)})q_{\ell}\left(\ell^{(k)}|m^{(k)},m^{(<k)}\right)/s^{(k)}$, where the $1/s^{(k)}$ factor arises from the determinant of the Jacobian of the change of variables from $(s^{(k)},r^{(k)})$ to $\ell^{(k)}$.
Algorithm 1 describes our procedure to evaluate the posterior $q_{\psi(x)}\left(\theta=(\tau,\ell)|x\right)$ of a phylogeny $\tau$ given set of sequences $x$. We use this procedure during the training phase to compute the loss function of Phyloformer 2, and at inference time when we want to evaluate the posterior probability of a phylogeny using a trained network. Of note, when evaluating $q_{\psi(x)}\left(\theta=(\tau,\ell)|x\right)$ the choice of one merge over several possibilities at each step is determined by the data, and does not depend one the network parameters. We therefore don’t need to differentiate through discrete operations during training even though the loss evaluation depends on a succession of discrete choices. By contrast, sampling from $q_{\psi(x)}\left(\theta=(\tau,\ell)|x\right)$ does require a sequence of discrete decisions that depend on the network parameters but we never need to differentiate through this process.
Sampling from the posterior given a set of sequences $x$ (Algorithm S.10) is very similar to the posterior evaluation procedure described above. The tree is iteratively built from EvoPF embeddings of $x$ by successively sampling merges from the softmin-parametrized conditional-merge probabilities. Similarly, branch-length at a step $k$ by sampling their sum from the shifted Gamma distribution and their ratio from the Beta distribution, whose parameters are obtained from EvoPF embeddings. Point estimation is also possible by using a greedy maximum a posteriori (MAP) approximation by using the mode of the estimated distributions, i.e., replacing softmin with argmin to always sample the most probable merge, and using the Gamma and Beta modes for branch lengths.
Of note, the resulting $q_{\psi}$ has limited expressivity for two reasons. First, the relative probabilities for merging each of the currently available pairs are computed based on the initial embeddings (i.e., the embeddings of those nodes that have not been chosen thus far are not updated as the recursion proceeds). In the true posterior, on the other hand, the relative posterior probability of being the next smallest cherry for a given pair of nodes in principle depends on previous merges. An update of the vector of embeddings at each step of the recursion could be implemented in a future version, but would be considerably more computationally intensive. Second, we model branch lengths posteriors with specific parametric distributions (Gamma and Beta) whereas the true posterior has no reason to match these analytical forms in general.
Algorithm 1 The BayesNJ Loss
function BayesNJ(Embeddings $\{\psi(x)\}=\{\mathbf{v}_{1},\ldots,\mathbf{v}_{N}\}$, merges $\{m^{(k)}\}$, branch lengths $\{\ell^{(k)}\}$)
   $\mathcal{L}_{\tau}\leftarrow 0,\mathcal{L}_{\ell}\leftarrow 0,\{c^{(1)}_{m}\}\leftarrow\{0\forall m\in\mathcal{S}_{0}^{2}\}$ Initialize log probabilities and constraints
   for all $k\in[1,\dots,N-2]$ do
    # Topological component
    $\{\text{score}_{ij}\}\leftarrow\left\{\text{SymmetricBilinear}(\mathbf{v}_{i},\mathbf{v}_{j})\right\}$ Score all pairs $(i,j)\in\mathcal{S}_{(k)}^{2}$
    $\{q_{m}(m=(i,j))\}\leftarrow\text{softmin}(\{\text{score}_{ij}\})$
    $\mathcal{L}_{\tau}\mathrel{+}=\log(q_{m}(m^{(k)})$
    # Branch length component
    $\mathbf{v}_{N+k}\leftarrow\text{SymmetricBilinear}(m^{(k)})$ Compute parent embedding
    $\left(\tilde{\alpha}^{(k)}_{\text{G}},\tilde{\lambda}^{(k)}_{\text{G}}\right)\leftarrow\text{SymmetricBilinear}(m^{(k)})$
    $\left(\alpha^{(k)}_{\text{G}},\lambda^{(k)}_{\text{G}}\right)\leftarrow 1+\text{softplus}\left(\tilde{\alpha}^{(k)}_{\text{G}},\tilde{\lambda}^{(k)}_{\text{G}}\right)+\varepsilon$
    $\left(\tilde{\alpha}^{(k)}_{\text{B}},\tilde{\beta}^{(k)}_{\text{B}}\right)\leftarrow\text{Bilinear}(m^{(k)})$
    $\tilde{\alpha}^{(k)}_{\text{B}}\leftarrow 1+\tilde{\alpha}^{(k)}_{\text{B}}$
    $\left(\alpha^{(k)}_{\text{B}},\beta^{(k)}_{\text{B}}\right)\leftarrow\text{softplus}\left(\tilde{\alpha}^{(k)}_{\text{B}},\tilde{\beta}^{(k)}_{\text{B}}\right)+\varepsilon$
    $\mathcal{L}_{\ell}\mathrel{+}=\log\text{PDF}_{\text{Gamma}}\left(\ell_{i}^{(k)}+\ell_{j}^{(k)}-c^{(k)}_{m^{(k)}}|\alpha^{(k)}_{\text{G}},\lambda^{(k)}_{\text{G}}\right)+\log\text{PDF}_{\text{Beta}}\left(\ell_{i}^{(k)}/\ell_{j}^{(k)}|\alpha^{(k)}_{\text{B}},\beta^{(k)}_{\text{B}}\right)$
    $\{c^{(k+1)}_{m}\}\leftarrow\left\{\max\left(c^{(k)}_{m},\ell_{i}^{(k)}+\ell_{j}^{(k)}\right),\forall m\in\mathcal{S}_{(k)}\right\}$ Update constraints
    # Prepare next iteration
    $\mathcal{S}_{(k+1)}=\left\{\mathcal{S}_{(k)}\cup u\right\}\setminus m^{(k)}$ Update mergeable nodes
   return $\mathcal{L}_{\tau},\mathcal{L}_{\ell}$
4 Experiments
4.1 Faster and more accurate point estimates of phylogenies
We trained Phyloformer 2 (PF2) over a large dataset simulated under similar priors to (Nesterenko et al., 2025). This dataset contains $\approx 1.3\cdot 10^{6}$ 50-taxa tree/MSA pairs simulated under a rescaled birth-death process—effectively corresponding to the prior $p(\theta)$—and the LG+G8 probabilistic model of evolution $p(x|\theta)$ (also see Section A.2.1).
The likelihood under this model is tractable, making it a favorable setting for maximum-likelihood methods—FastTree (Price et al., 2010) and IQTREE (Minh et al., 2020) in this experiment. We also include FastME (Lefort et al., 2015), a much faster but less accurate method that only requires to compute likelihoods of branches between pairs of sequences.
[…]
We then inferred greedy-MAP trees (see Section 3.3) on a test set of sequence alignments simulated under the same priors as the training set, with 10 to 200 taxa, obtained from Nesterenko et al. (2025). Figure 2a compares all tested methods using the normalized Robinson-Foulds distance between simulated and inferred tree—a classical metric to compare topologies (Robinson & Foulds, 1981). PF2 is a marked improvement over PF, with better topological accuracy across the whole range of tree/MSA size. For trees with 10 to 175 leaves, it also reconstructs trees with better accuracies than IQTree and FastTree, both state-of-the-art maximum-likelihood tree reconstruction methods. The edge of PF2 against maximum-likelihood methods working under the correct model likely arises from its estimating the posterior distribution using the correct prior—the same tree distribution is used to generate training and test samples—which reduces its variance without creating a bias.
[…]
Figure 2: (a) Topological performance for Phyloformer 2, measured by the normalized Robinson-Foulds distance. The alignments for which trees were inferred were taken from the original Phyloformer paper (Nesterenko et al., 2025) and were simulated under the LG+GC sequence model. (b) Runtime and (c) Memory usage for Phyloformer 2. The same GPU model as the original Phyloformer study was used to run Phyloformer 2 inference. The execution time for PF2 trained only on topology is also shown as a fainter version. Results for other methods are reported from the original Phyloformer paper (Nesterenko et al., 2025).
In order to disentangle the effects of the EvoPF module and the BayesNJ loss, we trained a PF2 model using an $\ell_{1}$ loss on pairwise distances, as was done with PF. After a similar number of training steps as PF, we inferred distance matrices from the PF test MSAs, and trees from these matrices using FastME (Lefort et al., 2015). $\text{PF2}_{\ell_{1}}$ yields trees with similar Kuhner-Felsenstein distances as PF, but a slightly better topological accuracies, especially for larger trees (See Figure S.2). This seems to indicate that although the EvoPF embedding scheme helps PF2 better predict topologies, most of the topological accuracy gain shown in Figure 2a is due to the BayesNJ loss.
4.2 Increased advantage under intractable probabilistic models of evolution
The main motivation for NPE is to allow well-specified inference under models whose likelihood is intractable. We assess the performance of PF2 in this setting on the same benchmark used in Nesterenko et al. (2025), that includes the Cherry (Prillo et al., 2023) model allowing for dependencies between evolution at distinct positions in the sequences and SelReg model (Duchemin et al., 2022) allowing for heterogeneous distributions between positions. Of note, a mistake in the generation of the Cherry dataset in Nesterenko et al. (2025) inflates the amount of dependency to an unrealistic level (see Appendix A.3) but we keep it as it still provides a proof of concept and to avoid the computational burden of re-training PF on a new dataset. Under both the Cherry and SelReg models, a fine-tuned PF2 model significantly outperforms the equivalent PF model in RF distance and is comparable to PF in terms of KF score (see Figure S.4). For both models and metrics, PF and PF2 outperform all other reference methods.
4.3 Estimating the phylogenetic posterior distribution
A major advance of PF2 compared to existing likelihood-free phylogenetic inference methods including PF is its ability to represent entire posterior distributions—as opposed to point estimates—over full phylogenies—as opposed to quartets. We assess the quality of this estimation by comparing against samples obtained from a long MCMC run of RevBayes (Höhna et al., 2016), a standard tool for Bayesian phylogenetic inference. We ran 10 parallel MCMC runs on a single 50 sequence alignment for $50,000$ iterations and $5,000$ burn-in iterations. We used a uniform prior on tree topologies and an Exponential distribution with $\lambda=10$ for branch lengths, and LG+G8 as the sequence evolution model. For a fairer comparison, we fine-tuned PF2 on tree/MSA pairs simulated under these priors before sampling from the posterior.
A common way to compare distribution of topologies is through the branches present in sampled trees. Every branch in a phylogeny defines a bipartition of the leaves, making them comparable across all possible trees and providing a softer metric than, e.g., the frequencies of full topologies. Figure S.3 shows that RevBayes produces a hard posterior distribution where most branches appear in either all or none of the sampled topologies. PF2 provides a softer posterior but a large agreement with RevBayes, as branches sampled in all RevBayes trees have a frequency larger than 0.6 in PF2, and those not sampled have a frequency mostly lower than 0.3. This is also consistent with our observation that the greedy MAP version of PF2 provides good point estimates, as sequentially sampling the highest probability merges is very likely to select the right branches. It is unclear whether the more diffuse distribution of bipartition probabilities in PF2 than in RevBayes is correct or not; it might be that on this particular data set RevBayes’s calibration is not optimal. Further investigation is necessary to evaluate the quality of the calibration of PF2’s bipartition probabilities.
5 Conclusion
[…]
One major limitation of Phyloformer 2 is its scalability, preventing its usage on more than 200 sequences. Future work should explore more efficient encoders (Wohlwend et al., 2025; Wang et al., 2025) or existing heuristics to build larger trees (Warnow, 2018; Jiang et al., 2024). In addition, Phyloformer 2 would currently produce poor estimates with no warning when presented with out of distribution inputs, in particular those far from its training data—a problem to which likelihood-based methods are immune as they do not require a learning phase. This issue is critical especially in cases where some areas have low probability under the prior, and could be mitigated by providing an assessment of the uncertainty of its prediction (Gal & Ghahramani, 2016; Lakshminarayanan et al., 2017).
For the sake of comparison with likelihood-based estimators, the present study mostly focused on tractable probabilistic models, but we expect Phyloformer 2 to reveal most of its potential by allowing inference under more realistic and complex scenarios (Latrille et al., 2021). Other promising directions are to handle unaligned sequences, and inference under broader probabilistic models of evolution embedding phylogenies such as population dynamics and co-evolution.
Acknowledgments
This work was funded by the Agence Nationale de la Recherche (ANR-20-CE45-0017). It was granted access to the HPC/AI resources of IDRIS under the allocation AD010714229R2 made by GENCI. The taxon silhouettes in Fig. 1 are modified from public domain images in the PhyloPic database. The authors also thank Vincent Garot and Evan Gorstein.
Appendix A Technical Appendices and Supplementary Material
Technical appendices with additional results, figures, graphs and proofs may be submitted with the paper submission before the full submission deadline (see above), or as a separate PDF in the ZIP file below before the supplementary material deadline. There is no page limit for the technical appendices.
A.1 Model architecture
Notation:
$i$, $j$ and $k$ denote sequence indices, $t$ and $l$ denote residue indices. Capitalized functions (e.g. Linear) are parametrized and learnable, whereas lowercase functions (e.g. sigmoid) are not.
The EvoPF module takes as input a multiple sequence alignments $\{\mathbf{x}_{it}\}$, where $\mathbf{x}_{it}$ is the one-hot encoded vector of the $t^{th}$ character of sequence $i$. The EvoPF module produces a set sequence embeddings $\{\mathbf{v}_{i}\}$ of fixed size $c_{s}=256$ for each sequence $i$, and a set a sequence pair embeddings $\{\mathbf{z}_{ij}\}$ of fixed size $c_{z}=512$ for each pair $(i,j)$ of sequences. Furthermore, the pair embedding vectors $\mathbf{z}_{ij}$ are projected to a single real number $z_{ij}$ to be used as a proxy for phylogenetic distance in the BayesNJ module.
Algorithm S.1 The EvoPF module
function EvoPF($N_{blocs}=12$, $\{\mathbf{x}_{it}\}$)
[…]
   for all $l\in[1,\ldots,N_{blocs}]$ do
[…]
   $z_{ij}=\text{Linear}(\mathbf{z}_{ij})$
   $\mathbf{v}_{i}=\text{mean}_{t}(\mathbf{v}_{it})$
   return $\{z_{ij}\},\{\mathbf{v}_{i}\}$ $\mathbf{z}_{ij}\in\mathbb{R}^{c_{z}}$, $\mathbf{v}_{i}\in\mathbb{R}^{c_{s}}$
A.1.1 Embedding modules
Algorithm S.2 MSA Embedding module
function MSAEmbedder($\{\mathbf{x}_{it}\}$, $c_{s}=128$)
   $\mathbf{a}_{it}=\text{Conv2D}(\mathbf{x}_{it})$ $a_{it}\in\mathbb{R}^{c_{s}}$
   $\mathbf{v}_{it}=\text{reLU}(\mathbf{a}_{it})$
   return $\{\mathbf{v}_{it}\}$
Algorithm S.3 Pair Embedding module
function PairEmbedder($\{\mathbf{x}_{it}\}$, $c_{z}=256$)
   $a_{it}=\text{reLU}(\text{Conv2D}(\mathbf{x}_{it})$) $a_{it}\in\mathbb{R}^{c_{z}}$
   $\mathbf{z}_{ij}=\text{mean}_{t}(a_{it}+a_{jt})$ $1\leq j<i\leq N$
   return $\{\mathbf{z}_{ij}\}$
A.1.2 The EvoPF MSA stack
Algorithm S.4 MSA Stack - Column-wise pair-biased gated self-attention
function MSAColAttentionWithPairBias($\{\mathbf{v}_{it}\},\{\mathbf{z}_{ij}\},N_{head}=4,c=c_{s}/N_{head}$)
   $\mathbf{v}_{it}\leftarrow\text{LayerNorm}(\mathbf{v}_{it})$
   $\mathbf{q}_{it}^{h},\mathbf{t}_{it}^{h},\mathbf{v}_{it}^{h},=\text{LinearNoBias}(\mathbf{v}_{it})$ $\mathbf{q}_{it}^{h},\mathbf{t}_{it}^{h},\mathbf{v}_{it}^{h}\in\mathbb{R}^{c},1\leq h\leq N_{head}$
   $b_{ij}^{h}=\text{LinearNoBias}(\mathbf{z}_{ij})$
   $\mathbf{g}_{it}^{h}=\text{sigmoid}(\text{Linear}(\mathbf{v}_{it})$) $\mathbf{g}_{it}\in\mathbb{R}^{c}$
   $a_{ijt}^{h}=\text{softmax}\left(\frac{1}{\sqrt{c}}\mathbf{q}_{it}^{h\top}\mathbf{t}_{jt}^{h}+b_{ij}^{h}\right)$
   $\mathbf{o}_{it}^{h}=\mathbf{g}_{it}^{h}\odot\sum_{j}a_{ijt}^{h}\mathbf{v}_{it}^{h}$
   $\mathbf{\tilde{s}}_{it}=\text{Linear}\left(\text{concat}_{h}(\mathbf{o}_{it}^{h})\right)$ $\mathbf{\tilde{s}}_{it}\in\mathbb{R}^{c_{s}}$
   return $\{\mathbf{\tilde{s}}_{it}\}$
Algorithm S.5 MSA Stack - Row-wise gated self-attention
function MSARowAttention($\{\mathbf{v}_{it}\},N_{head}=4,c=c_{s}/N_{head}$)
   $\mathbf{v}_{it}\leftarrow\text{LayerNorm}(\mathbf{v}_{it})$
   $\mathbf{q}_{it}^{h},\mathbf{t}_{it}^{h},\mathbf{v}_{it}^{h},=\text{LinearNoBias}(\mathbf{v}_{it})$ $\mathbf{q}_{it}^{h},\mathbf{t}_{it}^{h},\mathbf{v}_{it}^{h}\in\mathbb{R}^{c},1\leq h\leq N_{head}$
   $\mathbf{g}_{it}^{h}=\text{sigmoid}(\text{Linear}(\mathbf{v}_{it})$) $\mathbf{g}_{it}\in\mathbb{R}^{c}$
   $a_{itl}^{h}=\text{softmax}\left(\frac{1}{\sqrt{c}}\mathbf{q}_{it}^{h\top}\mathbf{t}_{il}^{h}\right)$
   $\mathbf{o}_{it}^{h}=\mathbf{g}_{it}^{h}\odot\sum_{l}a_{itl}^{h}\mathbf{v}_{il}^{h}$
   $\mathbf{\tilde{s}}_{it}=\text{Linear}\left(\text{concat}_{h}(\mathbf{o}_{it}^{h})\right)$ $\mathbf{\tilde{s}}_{it}\in\mathbb{R}^{c_{s}}$
   return $\{\mathbf{\tilde{s}}_{it}\}$
Algorithm S.6 MSA Stack - Transition
function MSATransition($\{\mathbf{v}_{it}\},n=4$)
   $\mathbf{v}_{it}\leftarrow\text{LayerNorm}(\mathbf{v}_{it})$
   $\mathbf{a}_{it}=\text{Linear}(\mathbf{v}_{it})$ $\mathbf{a}_{it}\in\mathbb{R}^{nc_{s}}$
   $\mathbf{v}_{it}\leftarrow\text{Linear}(\text{reLU}(\mathbf{a}_{it}))$
   return $\{\mathbf{v}_{it}\}$
Algorithm S.7 Communication - Outer product mean
function OuterProductMean($\{\mathbf{v}_{ik}\},c=32$)
   $\mathbf{v}_{it}\leftarrow\text{LayerNorm}(\mathbf{v}_{it})$
   $\mathbf{a}_{it},\mathbf{b}_{it}=\text{Linear}(\mathbf{v}_{it})$ $\mathbf{a}_{it},\mathbf{b}_{it}\in\mathbb{R}^{nc_{s}}$
   $\mathbf{o}_{ij}\leftarrow\text{flatten}(\text{mean}_{t}(\mathbf{a}_{it}\otimes\mathbf{b}_{it}))$ $\mathbf{o}_{ij}\in\mathbb{R}^{c^{2}}$
   $\mathbf{z}_{ij}=\text{Linear}(\mathbf{o}_{ij})$ $\mathbf{z}_{ij}\in\mathbb{R}^{c_{z}}$
   return $\{\mathbf{z}_{ij}\}$
A.1.3 The EvoPF pair stack
Algorithm S.8 Pair Stack - Gated self-attention
function PairAttention($\{\mathbf{z}_{ij}\},N_{head}=4,c=c_{z}/N_{head}$)
   $\mathbf{z}_{ij}\leftarrow\text{LayerNorm}(\mathbf{z}_{ij})$
   $\mathbf{q}_{ij}^{h},\mathbf{k}_{ij}^{h},\mathbf{v}_{ij}^{h},=\text{LinearNoBias}(\mathbf{z}_{ij})$ $\mathbf{q}_{ik}^{h},\mathbf{k}_{ij}^{h},\mathbf{v}_{ij}^{h}\in\mathbb{R}^{c},1\leq h\leq N_{head}$
   $\mathbf{g}_{ij}^{h}=\text{sigmoid}(\text{Linear}(\mathbf{z}_{ij})$) $\mathbf{g}_{ij}\in\mathbb{R}^{c}$
   $a_{ijk}^{h}=\text{softmax}\left(\frac{1}{\sqrt{c}}\mathbf{q}_{ij}^{h\top}\mathbf{k}_{jk}^{h}\right)$
   $\mathbf{o}_{ij}^{h}=\mathbf{g}_{ij}^{h}\odot\sum_{k}a_{ijk}^{h}\mathbf{v}_{ik}^{h}$
   $\mathbf{\tilde{z}}_{ij}=\text{Linear}\left(\text{concat}_{h}(\mathbf{o}_{ij}^{h})\right)$ $\mathbf{\tilde{z}}_{ik}\in\mathbb{R}^{c_{z}}$
   return $\{\mathbf{\tilde{z}}_{ij}\}$
Algorithm S.9 Pair Stack - Transition
function PairTransition($\{\mathbf{z}_{ij}\},n=4$)
   $\mathbf{z}_{ij}\leftarrow\text{LayerNorm}(\mathbf{z}_{ij})$
   $\mathbf{a}_{ij}=\text{Linear}(\mathbf{z}_{ij})$ $\mathbf{a}_{ij}\in\mathbb{R}^{nc_{z}}$
   $\mathbf{z}_{ij}\leftarrow\text{Linear}(\text{reLU}(\mathbf{a}_{ij}))$
   return $\{\mathbf{z}_{ij}\}$
A.1.4 Sampling from $q_{\psi(x)}(\theta=(\tau,\ell)|x)$ with BayesNJ
Algorithm S.10 Sampling from the posterior
For the Greedy-MAP approximation, we simply replace the softmin of scores with an argmin to sample merges, and replace samplings from Gamma and Beta distributions with the corresponding mode.
function SamplingBayesNJ(Embeddings $\{\psi(x)\}=\{\mathbf{v}_{1},\ldots,\mathbf{v}_{N}\}$)
   $\{c^{(1)}_{ij}\}\leftarrow\{0,\forall(i,j)\in[1,\dots,N-2]^{2}\}$ Initialize constraints
   for all $k\in[1,\dots,N-2]$ do
    # Topological component
    $\{\text{score}_{ij}\}\leftarrow\left\{\text{SymmetricBilinear}(\mathbf{v}_{i},\mathbf{v}_{j})\right\}$ Score all pairs $(i,j)\in\mathcal{S}_{(k)}^{2}$
    $m^{(k)}\sim\text{softmin}(\{\text{score}_{ij}\})$ Sample merge
    # Branch length component
    $\mathbf{v}_{N+k}\leftarrow\text{SymmetricBilinear}(m^{(k)})$ Compute parent embedding
    $\left(\tilde{\alpha}^{(k)}_{\text{G}},\tilde{\lambda}^{(k)}_{\text{G}}\right)\leftarrow\text{SymmetricBilinear}(m^{(k)})$
    $\left(\alpha^{(k)}_{\text{G}},\lambda^{(k)}_{\text{G}}\right)\leftarrow 1+\text{softplus}\left(\tilde{\alpha}^{(k)}_{\text{G}},\tilde{\lambda}^{(k)}_{\text{G}}\right)+\varepsilon$
    $\left(\tilde{\alpha}^{(k)}_{\text{B}},\tilde{\beta}^{(k)}_{\text{B}}\right)\leftarrow\text{Bilinear}(m^{(k)})$
    $\tilde{\alpha}^{(k)}_{\text{B}}\leftarrow 1+\tilde{\alpha}^{(k)}_{\text{B}}$
    $\left(\alpha^{(k)}_{\text{B}},\beta^{(k)}_{\text{B}}\right)\leftarrow\text{softplus}\left(\tilde{\alpha}^{(k)}_{\text{B}},\tilde{\beta}^{(k)}_{\text{B}}\right)+\varepsilon$
    $s^{(k)}\sim c^{(k)}_{m^{(k)}}+\text{Gamma}(\alpha^{(k)}_{\text{G}},\lambda^{(k)}_{\text{G}})$ Sample sum
    $r^{(k)}\sim\text{Beta}(\alpha^{(k)}_{\text{B}},\beta^{(k)}_{\text{B}})$ Sample ratio
    $\ell^{(k)}_{i}\leftarrow r^{(k)}\cdot s^{(k)}$
    $\ell^{(k)}_{j}\leftarrow s^{(k)}-\ell^{(k)}_{j}$
    # Prepare next iteration
    $\{c^{(k+1)}_{vw}\}\leftarrow\left\{\max\left(c^{(k)}_{vw},s^{(k)}\right),\forall(v,w)\in\mathcal{S}_{(k)}\right\}$ Update constraints
    $\mathcal{S}_{(k+1)}=\left\{\mathcal{S}_{(k)}\cup u\right\}\setminus m^{(k)}$ Update mergeable nodes
   return $\mathcal{L}_{\tau},\mathcal{L}_{\ell}$
A.2 Training runs
Model Starting weights Embedding dimensions ($c_{s}|c_{z}$) Training data Loss function Training time Batch size Scheduled epochs Target LR Warmup Selected step
[…]
(PF2) Random $(256|512)$ BD,LG+G8 BayesNJ 80h 9 30 $10^{-4}$ $10^{3}$ 89000
[…]
PF2 (PF2) $(256|512)$ BD,LG+G8,multi BayesNJ 26h 1-40 30 $10^{-6}$ $0.5\%$ 9276
[…]
Table S.1: Training run parameters. All runs were in a distributed data-parallel setting using 4 H100 GPUs. Final models are shown with their name in bold. More details on the training data in Section A.2.1 and Figure S.1.
Figure S.1: Training setting for all PF2 instances. PF2 instances are shown in rectangles, with a full outline and purple fill if they are final models used for inference in the main results, and a dotted outline and gray fill if they are used as the starting model for fine tuning runs. The loss functions used for each training or fine-tuning runs are shown in ovals. Finally the training (blue) and fine-tuning (yellow) datasets are also shown as cylindrical shapes. For each dataset, the number of training and validation examples is shown in the top section of the cylinder. The simulation priors are shown in the body of each cylinder : (1) the tree topology prior with the training tree size in parentheses, (2) the MSA evolutionary model. The tree priors are either (BD): rescaled Birth-death, described in (Nesterenko et al., 2025), or (U+Exp): Uniform tree topology with $\lambda=10$ exponentially distributed branch lengths available in RevBayes (Höhna et al., 2016). The number of figure in which the performance of a given model is studied is shown in the bottom right corner of the corresponding rectangle. Datasets are described in Section A.2.1
A.2.1 Training datasets
From (Nesterenko et al., 2025):
The 200k LG+GC, Cherry and Selreg datasets, used for training and fine-tuning models were obtained from the original paper. A full description of these datasets is available in the supplementary material of (Nesterenko et al., 2025)
LG+G8 dataset:
[…]
Multi-size LG+G8 dataset:
[…]
MCMC fine-tuning dataset:
This dataset was used to fine tune PF2 under the prior used in RevBayes (Höhna et al., 2016) to estimate the posterior distribution shown in Figure S.3. 72,007 training (resp. 1000 validation) trees were simulated with RevBayes with uniformly distributed topologies and branch lengths sampled from an exponential distribution with $\lambda=10$. MSAs were simulated along those trees using the same procedure as the main LG+G8 and the multi-size LG+G8 datasets described above.
A.3 The Cherry dataset in Nesterenko et al. (2025) contains unrealistic amounts of coevolution between sites
The Cherry dataset simulated in  Nesterenko et al. (2025) was simulated by using an incorrect rate matrix. As a result, the amount of co-evolution among pairs of coevolving residues was superior to what can be found in empirical data. The cause of this high amount of coevolution is a mistake in the use of the Cherry matrix Prillo et al. (2023). Instead of using the Cherry rate matrix to model pairs of interacting sites, the authors used the product of the rate matrix with its stationary frequencies. In standard models of molecular evolution, it is customary to represent a reversible substitution rate matrix $Q$ as a product of an exchangeability matrix $R$ and stationary frequencies $F$: $Q=R\times F$. The Cherry dataset in  Nesterenko et al. (2025) was actually simulated according to a matrix $Q^{\prime}=R\times F\times F=Q\times F$. The resulting data provides an example of data with extreme amounts of coevolution, which we use as an example of strong departure from standard models of sequence evolution as implemented in e.g., IQ-Tree Minh et al. (2020).
A.4 Supplementary Figures
Figure S.2: How does PF2 trained with an L1 loss compare to PF? PF2 was trained under the same conditions as the original PF on LGGC trees with an L1 loss on pairwise phylogenetic distances. The Robinson-Foulds distance (left) shows the topological reconstruction accuracy, while the Kuhner-Felsenstein distance (right) takes both topology and branch-lengths into account.
Figure S.3: Comparison of split frequencies over samples from the posterior of a single 50 sequences MSA, between RevBayes MCMC (x-axis) and PF2 (y-axis). The orange cross marker indicates splits that are present in the tree along which the MSA used for sampling has been simulated.
Figure S.4: Average PF2 performance under intractable-likelihood models, measured on 50-tip trees. The PF and PF2 versions are fine tuned to either Cherry (top row) or the SelReg (bottom row) data. Error bar show the 95% CI computed with 1000 bootstrap samples.
Figure S.5: Topological performance for non fine-tuned topology-only Phyloformer 2, measured by the normalized Robinson-Foulds distance. The MSA dataset and compared method results are the same as Figure 2
Figure S.6: PF2 memory usage (in GB) scaling w.r.t number of sequences and sequence length measured on an H100 GPU